OpenStudy (anonymous):

PLEASE HELP Write cos(t) in terms of csc(t) if the terminal point determined by t is in quadrant II.

OpenStudy (amistre64):

csc is in inverted sin; and cos is a shifted sin...

OpenStudy (amistre64):

cos(t) = sin(t+ (pi/2)) csc(t) = 1/sin(t) sooooo..... cos(t) = 1/csc(t+ (pi/2))

OpenStudy (anonymous):

My only options are A. -((csc t)/(sqrt 1 + csc^2t)) B. -((1)/(sqrt 1 + csc^2t)) C. - ((sqrt csc^2t -1)/(csc t)) D. ((1)/(sqrt csc^2t -1)) E. -((1- csc^2t)/(csc t))

OpenStudy (amistre64):

csc is negative in q2; so d is out

OpenStudy (amistre64):

we can work it backwards as well; pick an angle you know the cos and csc to in Q2 and we what fits :)

OpenStudy (amistre64):

cos(120) = 1/2 ; the sin(120) = -sqrt(3)/2; csc(120) = -2/sqrt(3)

OpenStudy (amistre64):

-2 1 --------*k = --- sqrt(3) 2 sqrt(3) k = - -------- 4

OpenStudy (amistre64):

i cant get a clear picture in my head of what the problem is asking for yet....

OpenStudy (anonymous):

csc(t) = 1/sin(t) \[cos(t) = \sqrt{1-\sin^2t}\] \[\cos(t) = \sqrt{1-1/\csc^2t}\]

OpenStudy (anonymous):

It is in the II quadrant, so C is the option!